OpenStudy (anonymous):

Need verification: Apply newton's method with initial approximation x1=-4 to find the 2nd and 3rd approximations x2 and x3 to the real root of the equation

OpenStudy (anonymous):

\[x^3 -9x +28 = 0\] my answer is -4 for both x2 and x3.. can anyone verify this?

OpenStudy (ksaimouli):

can u explain ur steps

OpenStudy (anonymous):

\[x2 = x1 - \frac{ f(x1) }{ f'(x1) }\] \[x3 = x2 - \frac{ f(x2) }{ f'(x2) }\]

OpenStudy (ksaimouli):

i am getting 0

OpenStudy (ksaimouli):

x1

OpenStudy (anonymous):

\[x2 = (-4) - (\frac{ (-4)^3 -9(-4) +28 }{ 3(-4) - 9 })\]

OpenStudy (anonymous):

which is equal to -4, and repeats for x3

OpenStudy (anonymous):

wait there is a power next to the -4 on the bottom

OpenStudy (anonymous):

i just forgot that, but i still get -4 for x2

OpenStudy (anonymous):

so its (-4) - (0/39)....right?

OpenStudy (ksaimouli):

hmm but i got x1=0 so should'nt the x2=0-F(0)/(f'(0))

OpenStudy (anonymous):

x1 is already given to be x1 = -4

OpenStudy (anonymous):

or is initial approximation x0?

OpenStudy (ksaimouli):

ohh yup then ur right

OpenStudy (anonymous):

im confused though..so is initial approximation (-4) x0 or x1?

OpenStudy (anonymous):

i know it says x1 in the question, but im not sure

OpenStudy (anonymous):

i think it is probably x1

OpenStudy (ksaimouli):

it is x1 clearly stated in question

OpenStudy (anonymous):

do you get -4 for both x2 and x3 now when you try it?

OpenStudy (ksaimouli):

yup

OpenStudy (anonymous):

ok.. thanks